# Can LLM-Augmented autonomous agents cooperate?, An evaluation of their cooperative capabilities through Melting Pot

### v1

@@ Appendix A Descriptions generated for the objects in the environment
In Table 1 we show all the natural language descriptions generated to represent the relevant objects and events of the Commons Harvest scenario of Melting Pot.
Object/Event |
Description |
@@
Observed agent |
@@
Observed grass to grow apples at position |
@@
Observed an apple at position |
@@
Observed tree |
@@
Someone was attacked at position |
@@
Observed a ray beam from an attack at position |
@@
Observed that agent |
@@
Observed that the grass at position |
@@
Observed that grass to grow apples appeared at position |
@@
Observed that an apple grew at position |
@@
There are no observations: You were attacked by agent |
The agent is out of the game |
There are no observations: you’re out of the game. |
[…]
@@ Appendix B Knowledge about the world given to agents
The Listing 1 shows the raw world description passed to the agents. This is the only information agents have about the environment.
@@ Appendix C React prompt
The Listing 2 shows the whole prompt used in the react module. This prompt was used to let the agent decide whether to react or not to the current observations, where react means to change the plan and generate a new action. The inputs that this prompt receives are the following in order: name, world context, current observations, current plan, actions to take if any, changes observed in the game state, game time, and agent’s personality,
@@ Appendix D Plan prompt
The Listing 3 shows the raw prompt used in the plan module. This prompt helps the agent make a high-level plan and define several goals to guide its actions. The inputs that this prompt receives are the following in order: name, world context, current observations, current plan, reflections, reason to react, agent’s personality, and changes observed in the game state.
@@ Appendix E Reflection prompts
The Listing 4 shows the raw prompt used in the first part of the reflections module i.e. question formulation. The inputs for this prompt are the following: name, world context, accumulated observations since the last reflection, and agent’s personality.
The prompt used in the insight generation part that takes place in the reflect module is shown in Listing 5, its corresponding inputs are the following: name, world context, group of memories retrieved for each generated question in the first part, and agent’s personality.
@@ Appendix F Act prompt
The Listing 6 shows the raw prompt used in the act module. This prompt is in charge of deciding which action to take. The inputs that this prompt receives are the following: name, world context, current plan, the most recent ten reflections, current observations, number of actions to generate, set of valid actions, current goals, agent’s personality, position of the known trees, portion of the map explored, previous actions, and changes observed in the game state.
@@ Appendix G Simulations cost
| Experiment | Avg. Simulation cost | Avg. execution time (minutes) |
@@
| Set 1 - All Coop with def111For this experiment the models of OpenAI were used through Azure. | ||
@@
| Set 2 - Agents versus Bots | ||
| Set 3 - All aware one selfish | ||
[…]

### v2

@@ Appendix A Descriptions generated for the objects in the environment
In Table 1, we show all the natural language descriptions generated to represent the relevant objects and events of the Commons Harvest scenario of Melting Pot.
| Object/Event | Description |
@@
Observed agent <agent_name> at position [<x>, <y>].
@@
Observed grass to grow apples at position [<x>, <y>]. This grass belongs to tree <tree_id>.
@@
Observed an apple at position [<x>, <y>]. This apple belongs to tree <tree_id>.
@@
Observed tree <tree_id> at position [<x>, <y>]. This tree has apples_number apples remaining and grass_number grass for apples growing on the observed map. The tree might have more apples and grass on the global map.
@@
Someone was attacked at position [<x>, <y>].
@@
Observed a ray beam from an attack at position [<x>, <y>].
@@
Observed that agent agent_name took an apple from position [<x>, <y>].
@@
Observed that the grass at position [<x>, <y>] disappeared.
@@
Observed that grass to grow apples appeared at position [<x>, <y>].
@@
Observed that an apple grew at position [<x>, <y>].
@@
There are no observations: You were attacked by agent agent_name and currently you’re out of the game.
| The agent is out of the game | There are no observations: you’re out of the game. |
@@ Appendix B Knowledge about the world given to agents
The Listing 2 shows the raw world description passed to the agents. This is the only information agents have about the environment.
@@ Appendix C React Prompt
The Listing 3 shows the entire prompt used in the react module. This prompt enables the agent to decide whether to react to the current observations—where reacting implies altering the plan and generating a new action. The prompt receives inputs in the following order: name, world context, current observations, current plan, actions to take if any, changes observed in the game state, game time, and agent’s personality.
@@ Appendix D Plan prompt
The Listing 4 shows the raw prompt used in the plan module. This prompt helps the agent make a high-level plan and define several goals to guide its actions. The inputs that this prompt receives are the following in order: name, world context, current observations, current plan, reflections, reason to react, agent’s personality, and changes observed in the game state.
@@ Appendix E Reflection prompts
The Listing 5 shows the raw prompt used in the first part of the reflections module i.e. question formulation. The inputs for this prompt are the following: name, world context, accumulated observations since the last reflection, and agent’s personality.
The prompt used in the insight generation part that takes place in the reflect module is shown in Listing 6, its corresponding inputs are the following: name, world context, group of memories retrieved for each generated question in the first part, and agent’s personality.
@@ Appendix F Act prompt
The Listing 7 shows the raw prompt used in the act module. This prompt is in charge of deciding which action to take. The inputs that this prompt receives are the following: name, world context, current plan, the most recent ten reflections, current observations, number of actions to generate, set of valid actions, current goals, agent’s personality, position of the known trees, portion of the map explored, previous actions, and changes observed in the game state.
@@ Appendix G Simulations Cost
| Avg. Simulation | Avg. Execution | |
| Experiment | Cost ($) | Time (minutes) |
@@
| Set 1 - All Coop with def111For this experiment, the models of OpenAir were used through Classic. | ||
@@
| Set 2 - Agents vs. Bots | ||
| Set 3 - All aware one selfish |